# Vision Expert-guided Visual Tokenizer and Manufacturing Process

### v1

@@ 1 Introduction
Figure 1: Main workflow of Triad. As shown in the figure, the query image is first passed through EG-RoI with the suspicious regions predicted by the vision expert. These regions are cropped keeping a higher resolution than the query image and then encoded along with it. The text input includes a basic prompt, the manufacturing process for the specific product (this kind of cable for this case) and a question about anomaly detection. In this case, Triad finds out that the issue is the wrong insulation color, as the manufacturing process mentioned the cable is composed of three different color wires.
@@
Industrial Anomaly Detection (IAD) plays a crucial role in modern manufacturing, where the continuous iteration of the product design and diverse inspection criteria call for robust and adaptable inspection solutions. Although large multimodal models (LMMs) have recently shown remarkable success on general-purpose tasks, their performance in IAD remains noticeably below expectations. We attribute this discrepancy to two key factors.
@@
First, current LMMs are typically trained to align vision and language modalities for broad semantic understanding, yet they lack the specialized ability to identify and focus on potential defect regions. Industrial defects often appear subtly against complex backgrounds or in close proximity to functional components, making a purely global, coarse alignment insufficient. To address this challenge, we propose an expert-guided region-of-interest tokenizer (EG-RoI). Inspired by AnyRes module in LLaVA [23, 22, 16], EG-RoI extracts high-resolution and potentially anomalous regions of interest identified by existing IAD methods. By explicitly highlighting anomalous and normal regions during training, we enable the model to learn targeted visual cues indicative of attributes of products and defects. During inference, these high resolution suspicious region predicted by vision experts further serve as key references for more precise anomaly detection.
@@
Second, prevailing IAD approaches treat defects as isolated visual phenomena, neglecting their intrinsic relationship to manufacturing workflows. In reality, anomalies arise from process deviations (material impurities, assembly errors, or equipment malfunctions) that propagate through production stages. Existing methods, which rely on image-level comparisons or defect taxonomies, fail to leverage this causal knowledge. To bridge this gap, we introduce a manufacturing-driven industry anomaly detection paradigm that weaves relevant manufacturing processes into the reasoning process. Our framework incorporates: (1) An instruction-tuning dataset with attribute-rich captions, InstructIAD, spanning a variety of products and defect types from existing IAD datasets; (2) Chain-of-Thought with Manufacturing (CoT-M), a data organization strategy that synthesizes defect causality by linking anomalies to specific manufacturing steps. CoT-M expands InstructIAD by (i) editing normal product captions at the attribute level to simulate product evolution and defect occurrence and (ii) generating Chain-of-Thought processes grounded in manufacturing steps via GPT. When captions are unavailable, CoT-M employs a checklist-style template mirroring human inspector workflows. Ultimately, we propose Triad, a novel LMM-based method incorporating an expert-guided region-of-interest tokenizer and manufacturing process for industrial anomaly detection. We evaluate Triad on three benchmarks, MVTec-AD [2], WFDD [6], and PCB-Bank [33], and assess results under 0-/1-shot settings. Experiments show that Triad-ov-7B not only achieves competitive performance against both general-purpose and domain-specific LMMs but also demonstrates further improvements when accounting for manufacturing processes. Notably, with a single reference image (1-shot), Triad achieves 94.1% on MVTec-AD, surpassing Qwen2-VL-72B [31] by 2.1%. Extensive analyses show the ability of Triad to comprehend complex manufacturing workflows and extend to unseen processes.
@@ 2.2 Large Multimodal Models for IAD
Recent progress in LMMs has led to significant advances across a variety of vision-language tasks, driven by improvements in visual-encoder architectures such as Q-Former [17], adaptive visual encoding [34], and spatial-aware visual sampling [38]. Several IAD-centric methods attempt to incorporate such models by supplying textual guidance. For instance, Customizable-VLM [32] tailors off-the-shelf LMMs to anomaly detection using class-specific prompts and normality descriptions, while MMAD [13] employs a defect-aware strategy to integrate a database of normal product features and defect descriptions. Although these textual cues enhance performance, they still fall short of capturing the underlying interactions among manufacturing steps that directly lead to unforeseen defects.
The most closely related efforts include AnomalyGPT [10] and Myriad [19], which fine-tunes LMMs to the IAD domain. AnomalyGPT uses an image decoder to predict pixel-level anomaly maps and then applies a prompt learner to guide large language models, but its reliance on coarse-level image tokens and predicted masks can lead to loss of fine-grained details. Myriad leverages pretrained IAD models as vision experts to resample fine-grained features via Q-Former [17], effectively transferring general-purpose visual perception into the IAD domain. However, their reliance on simulated anomaly data and limited vision-language alignment undermines their ability to capture fine-grained visual attributes and understand manufacturing processes, ultimately leading to suboptimal performance. To bridge these gaps, our work, Triad, first enhances attribute perception by introducing the CoT-M and EG-RoI to further enable manufacturing-assisted anomaly detection.
@@ 3.1 Problem definition
While standard IAD tasks (e.g., [10, 19]) primarily focus on visual cues, manufacturing-driven IAD leverages intricate manufacturing process details, such as multi-step production procedures, to support a deeper analysis. This paradigm demands accurate visual attribute recognition and sophisticated reasoning that accounts for the interplay between the visual cues and the contextual manufacturing information, enabling the more precious detection of defects arising from complex process variations.
@@ 3.2 Training Objectives
Based on the problem defined before, our input, noted as $X$, is composed of vision part $X_{V}$ and language part $X_{L}$, with language annotation $Y$ having $n$ tokens $(y_{1},y_{2},...,y_{n})$ from our training dataset including InstructIAD and CoT-M. $X_{V}$ is composed of query image $X_{img}$ and extra patches produced by EG-RoI.
@@ 3.4 Chain-of-Thought with Manufacturing
To achieve manufacturing-aware reasoning, we introduce CoT-M: a data organization strategy that infuses Chain-of-Thought (CoT) reasoning with manufacturing processes. CoT-M is built upon InstructIAD, and is structured to enhance anomaly detection by integrating product attributes and manufacturing-related information.
CoT-M extends InstructIAD by (i) diversifying product and defect descriptions and (ii) synthesizing detailed manufacturing-aware reasoning trajectories. The overall pipeline, illustrated in Fig. 2, starts with the attribute-level captions in InstructIAD, which then serve as inputs to GPT. GPT generates reasoning sequences by combining the image caption with manufacturing processes to produce visual CoT answer pairs (image, manufacturing process, reasoning process). When caption information is unavailable (e.g., in MVTec-LOCO AD’s “juice” category), we apply a CoT template designed to mimic a human inspector’s checklist. This template cross-references each coarse label with a LLM-generated checklist from the manufacturing process to determine which visible attribute is defective. The final checklist thus serves as the chain-of-thought explanation.
To simulate product evolution and defect occurrence, we further augment the data using LLMs. Beginning with a normal product description in InstructIAD, we modify product attributes (e.g., color, component types, quantity) and inject defects at randomized locations. This process yields diverse color variations, component changes, and defect patterns. After generating the accompanying reasoning trajectories with the augmented descriptions and manufacturing processes, we manually filter out any hallucinated or erroneous samples. The final output including an augmented caption, corresponding manufacturing processes, and generated reasoning process forms a new textual CoT answer pair. By combining InstructIAD and samples organized by CoT-M, our goal is to strengthen the model’s ability to enhance anomaly detection while incorporating a deeper understanding of manufacturing workflows.
@@ 3.5 Expert-guided RoI module
In LMMs, the image is encoded into tokens by a visual encoder such as CLIP [27] and SigLIP [37], thus the encoder resizes the resolution of the input images to a fixed size. To offset the effect of limited resolution, LLaVA-Next [22] has proposed a technique called AnyRes by entering extra large patches of the original image. This technique improves performance and has been further developed in LLaVA-OneVision [16]. However, dividing the whole image into patches as extra patches is memory consuming and redundant for IAD as most patches of a defective image do not include defects. Thus, our EG-RoI module, shown in Fig. 3, introduces vision experts to provide more precise regions to further improve the resolution and reduce redundant tokens. For vision experts, we use the traditional zero-shot IAD method such as MuSc [18], AnomalyClip [40] and April-GAN [7]. The anomaly maps predicted by vision experts are normalized and binarized to select pixels with scores above 0.9 as suspicious regions. The resolution of these regions will be kept as much as possible by choosing a fixed-size bounding box and cropping them from the original image. Due to the context length limit, we merge boxes overlapped too much and reduce the number of boxes to 4. The original image is resized to fit the size of the visual encoder as the base image. Then the extra patches of suspicious regions are encoded by the image encoder along with the base image. After being encoded as tokens, these extra patches are applied with average pooling to reduce memory cost and flattened before concatenating with the base image features. The last operation embeds them into text tokens and inputs the whole sequence of tokens into the language model. Though guided by visual-experts, these suspicious regions still need inspection by LMM. With EG-RoI, suspicious local regions are provided in a relatively high resolution to LMM for more detailed inspection.
@@ 4 Experiments
GPT-4o [11] - 82.2% 67.9% (14.3%) $\downarrow$ 78.5% 77.3% (1.2%) $\downarrow$
Qwen2-VL [31] 2B 77.0% 46.7%(30.3%) $\downarrow$ 70.6% 45.2%(25.4%) $\downarrow$
LLava-1.6 [22] 7B 76.9% 75.9% (1.0%) $\downarrow$ 63.8% 64.0% (0.2%) $\downarrow$
MiniCPM-V [34] 8B 62.3% 51.6%(10.7%) $\downarrow$ 70.3% 52.1%(18.2%) $\downarrow$
LLaVA-OneVision-si [16] 7B 77.7% 60.6%(17.1%) $\downarrow$ 65.2% 61.4% (3.8%) $\downarrow$
LLaVA-OneVision-ov [16] 7B 91.0% 80.8%(10.2%) $\downarrow$ 79.8% 80.3% (0.5%) $\uparrow$
Qwen2-VL [31] 7B 84.4% 61.1%(23.3%) $\downarrow$ 74.4% 61.4%(13.0%) $\downarrow$
Qwen2-VL [31] 72B 87.1% 79.5% (7.6%) $\downarrow$ 81.1% 74.2% (6.9%) $\downarrow$
LLaVA-OneVision-ov [16] 72B 87.3% 75.5%(11.8%) $\downarrow$ 75.0% 74.6% (0.4%) $\downarrow$
Myriad [19] 7B 79.3% 81.5% (2.5%) $\uparrow$ 60.5% 61.7% (1.2%) $\uparrow$
Triad-llava-1.6 7B 85.0% 87.5% (2.5%) $\uparrow$ 67.3% 69.9% (2.6%) $\uparrow$
Triad-ov 7B 91.2% 92.6% (1.4%) $\uparrow$ 80.2% 81.1% (0.9%) $\uparrow$
Our experiments mainly focus on how industrial knowledge boosts LMMs in IAD tasks through our method, including common results from general LMMs and LMM-based IAD methods and ablation experiments of different industrial contexts.
@@
To preserve the generalization of the base models, we supplement our IAD-related instruction data by sampling 12K pairs from the original fine-tuning datasets of both LLaVA-1.6 and LLaVA-OneVision. Moreover, since LLaVA-1.6 lacks inherent multi-image processing capabilities, we constructed a simple dual-image caption dataset using the COCO subset [21] from the ShareGPT4V dataset [5]. This dataset is exclusively used in the 1-shot setting to provide basic multi-image support. Given the distinct objectives of 0-/1-shot anomaly detection, we offer separate versions of Triad for each setting, with data organized by CoT-M with specific instruction. For the 1-shot version of Triad-llava-1.6, we integrate the zero-shot model with the one-shot model using a Confidence Voting Mechanism (see Sec. E.1) to mitigate its multi-image inability.
@@
For zero-shot evaluation, the instruction follows the same formation as the anomaly detection task shown in Fig. 2. For one-shot evaluation, general LMMs that support multi-image directly input the query and a normal reference image with a similar prompt with zero-shot, only the question is replaced with:
@@
LMM-based IAD method AnomalyGPT [10] and Myriad [19] are tested according to their own instructions. Because of their low instruction following ability, we retrieve the keyword yes or no for accuracy calculation. Detailed demonstration of the retrieve method and detailed sub-dataset results can be found in Sec. B.
Baseline In this study, we perform extensive evaluations against several robust baselines, including AnomalyGPT [10] and Myriad [19]. General-purpose LMMs exhibit enhanced reasoning capabilities as a result of being fine-tuned on complex reasoning data spanning diverse domains, including mathematics, chart interpretation, and document analysis. To evaluate their performance, we benchmark our approach against state-of-the-art general-purpose LMMs, including Qwen2-VL [31], MiniCPM-V [34], the LLaVA series [22, 16], and the closed-source GPT-4o [11].
@@ 4.1 Main results
Table 2 shows 1-shot results. Providing a single reference image generally degrades the performance of most general-purpose LMMs due to their limited instruction-following capabilities. Two exceptions, LLaVA-OneVision-ov-7B (91.5% vs. 91.0% in 0-shot) and Qwen2-VL-72B (92.0% vs. 87.3% in 0-shot), still cannot fully utilize manufacturing information for anomaly detection (accuracy down by 5.7% and 1.6%, respectively). Meanwhile, Triad-llava-1.6-7B achieves comparable performance to both AnomalyGPT and Myriad under 1-shot conditions, and significantly outperforms them with manufacturing processes (88.4% vs. 85.4% for Myriad and 80.9% for AnomalyGPT). Leveraging LLaVA-OneVision-ov-8B—a carefully tuned multi-image baseline—Triad-ov-8B achieves 92.9% on MVTec-AD in the 1-shot setting, further improving to 94.1% when incorporating manufacturing processes. These results suggest that even a strong base model can benefit from the Manufacturing processes. Overall, our findings demonstrate that Triad offers robust and flexible anomaly detection across both zero-shot and few-shot scenarios, while the manufacturing-driven IAD paradigm consistently delivers strong performance gains across diverse IAD settings.
@@
GPT-4o [11] - 77.6% 72.5% (5.1%) $\downarrow$
Qwen2-VL [31] 2B 32.1% 30.7% (1.4%) $\downarrow$
LLava-1.6 [22] 7B 72.3% 76.0% (3.7%) $\uparrow$
LLaVA-OneVision-ov [16] 7B 91.5% 85.2% (5.7%) $\downarrow$
Qwen2-VL [31] 7B 81.7% 84.5% (2.8%) $\uparrow$
Qwen2-VL [31] 72B 92.0% 90.4% (1.6%) $\downarrow$
AnomalyGPT [^1] AnomalyGPT adds hints mentioning defect types for each product in MVTec-AD test set. [10] 7B 86.1% 80.9% (5.2%) $\downarrow$
Myriad [19] 7B 87.4% 85.4% (2.0%) $\downarrow$
Triad-llava-1.6 7B 87.7% 88.4% (0.7%) $\uparrow$
Triad-ov 7B 92.9% 94.1% (1.2%) $\uparrow$
@@
Figure 5: Qualitative evaluation of Triad-llava-ov-7B on MVTec-AD [2]. MuSc [18] is used as experts. The related manufacturing process is shown in (a). Three representative examples from the “Pill” class are shown: (a) a normal pill exhibits red speckles on its surface which—although typically interpreted as defects—are a natural outcome of the manufacturing process. (b) Pills with unintended additive variations indicate defect. (c) a case where Triad correctly predicts a defect despite errors in the expert outputs.
@@
Figure 5 presents three representative cases. In the first example (Fig. 5(a)), a surface discoloration might typically be viewed as a defect (as GPT-4o incorrectly predicts); however, Triad recognizes that the discoloration corresponds to the cranberry additive described in the manufacturing process, leading to a correct classification. Thanks to the EG-RoI module’s ability to capture fine-grained visual cues and the data organization via CoT-M, Triad confidently aligns the observed appearance with the relevant manufacturing steps.
In the second example (Fig. 5(b)), a pill containing “yellow speckles” is mistakenly introduced into a production pipeline that should only produce pills with “red speckles.” Triad correctly identifies the inconsistency between the visual observation and the manufacturing procedures. GPT-4o, though aware that red speckles are the standard, fails to accurately match the color of the speckles in question. This underscores the necessity of precise attribute recognition for reasoning.
Finally, the third example (Fig. 5(c)) highlights Triad’s resilience to vision expert errors. In this case, the expert fails to detect a scratch on the surface. Triad still identifies the scratch independently and also notes a yellow discoloration, demonstrating its capacity to rectify external region proposal mistakes by integrating visual analysis with contextual defect reasoning.

### v2

# Expert-guided Region-of-Interest Tokenizer and Manufacturing Process

@@ 1 Introduction
Figure 1: Main workflow of Triad. As shown in the figure, the query image is first passed through EG-RoI with the suspicious regions predicted by the vision expert. These regions are cropped to keep a higher resolution than the query image and then encoded along with it. The text input includes a basic prompt, the manufacturing process for the specific product (this kind of cable for this case) and a question about anomaly detection. In this case, Triad finds out that the issue is the wrong insulation color, as the manufacturing process mentioned that the cable is composed of three different color wires.
@@
Industrial Anomaly Detection (IAD) plays a crucial role in modern manufacturing, where continuous iteration of product design and diverse inspection criteria call for robust and adaptable inspection solutions. Although large multimodal models (LMMs) have recently shown remarkable success on general-purpose tasks, their performance in IAD remains noticeably below expectations. We attribute this discrepancy to two key factors.
@@
First, current LMMs are typically trained to align vision and language modalities for broad semantic understanding, yet they lack the specialized ability to identify and focus on potential defect regions. Industrial defects often appear subtly against complex backgrounds or in close proximity to functional components, making a purely global, coarse alignment insufficient. To address this challenge, we propose an expert-guided region-of-interest tokenizer (EG-RoI). Inspired by AnyRes module in LLaVA [23, 22, 16], EG-RoI extracts high-resolution and potentially anomalous regions of interest identified by existing IAD methods. By explicitly highlighting anomalous and normal regions during training, we enable the model to learn targeted visual cues indicative of attributes of products and defects. During inference, the high resolution suspicious regions predicted by vision experts further serve as key references for more precise anomaly detection.
@@
Second, prevailing IAD approaches treat defects as isolated visual phenomena, neglecting their intrinsic relationship to manufacturing (MFG) workflows. In reality, anomalies arise from process deviations (material impurities, assembly errors, or equipment malfunctions) that propagate through production stages. Existing methods, which rely on image-level comparisons or defect taxonomies, fail to leverage this causal knowledge. To bridge this gap, we introduce a manufacturing-driven industry anomaly detection paradigm that weaves relevant manufacturing processes into the reasoning process. Our framework incorporates: (1) An instruction-tuning dataset with attribute-rich captions, InstructIAD, spanning a variety of products and defect types from existing IAD datasets; (2) Chain-of-Thought with Manufacturing (CoT-M), a data organization strategy that synthesizes defect causality by linking anomalies to specific manufacturing steps. CoT-M expands InstructIAD by (i) editing normal product captions at the attribute level to simulate product evolution and defect occurrence and (ii) generating Chain-of-Thought processes grounded in manufacturing steps via GPT. When captions are unavailable, CoT-M employs a checklist-style template mirroring human inspector workflows. Ultimately, we propose Triad, a novel LMM-based method incorporating an expert-guided region-of-interest tokenizer and manufacturing process for industrial anomaly detection. We evaluate Triad on three benchmarks, MVTec-AD [2], WFDD [6], and PCB-Bank [33], and assess results under 0-/1-shot settings. Experiments show that Triad-ov-7B not only achieves competitive performance against both general-purpose and domain-specific LMMs but also demonstrates further improvements when accounting for manufacturing processes. Notably, with a single reference image (1-shot), Triad-ov-7B achieves 94.1% on MVTec-AD, surpassing Qwen2-VL-72B [31] by 2.1%. Extensive analyses show the ability of Triad to comprehend complex manufacturing workflows and extend to unseen processes.
@@ 2.2 Large Multimodal Models for IAD
Recent progress in LMMs has led to significant advances across a variety of vision-language tasks, driven by improvements in visual-encoder architectures such as Q-Former [17], adaptive visual encoding [34], and spatial-aware visual sampling [38]. Several IAD-centric methods attempt to incorporate such models by supplying textual guidance. For instance, Customizable-VLM [32] tailors off-the-shelf LMMs to anomaly detection using class-specific prompts and normality descriptions, while MMAD [13] employs a defect-aware strategy to integrate a database of normal product features and defect descriptions. Although these textual cues enhance performance, they still fall short of capturing the underlying interactions among manufacturing steps that directly lead to defects.
The most closely related efforts include AnomalyGPT [10] and Myriad [19], which fine-tunes LMMs to the IAD domain. AnomalyGPT uses an image decoder to predict pixel-level anomaly maps and then applies a prompt learner to guide large language models, but its reliance on coarse-level image tokens and predicted masks can lead to loss of fine-grained details. Myriad leverages pretrained IAD models as vision experts to resample fine-grained features via Q-Former [17], effectively transferring general-purpose visual perception into the IAD domain. However, their reliance on simulated anomaly data and limited vision-language alignment undermines their ability to capture fine-grained visual attributes and understand manufacturing processes, ultimately leading to suboptimal performance. To bridge these gaps, our work, Triad, first enhances attribute perception by introducing the CoT-M and EG-RoI, allowing manufacturing-driven anomaly detection.
@@ 3.1 Problem definition
While standard IAD tasks (e.g., [10, 19]) primarily focus on visual cues, manufacturing-driven IAD leverages intricate manufacturing process details, such as multi-step production procedures, to support a deeper analysis. This paradigm demands accurate visual attribute recognition and sophisticated reasoning that accounts for the interplay between visual cues and contextual manufacturing information, allowing more precise detection of defects arising from complex process deviations.
@@ 3.2 Training Objectives
Based on the problem defined before, our input, noted as $X$, is composed of vision part $X_{V}$ and language part $X_{L}$, with language annotation $Y$ having $n$ tokens $(y_{1},y_{2},...,y_{n})$ from our training dataset, including InstructIAD and CoT-M. $X_{V}$ is composed of query image $X_{img}$ and extra patches produced by EG-RoI.
@@ 3.4 Chain-of-Thought with Manufacturing
To achieve manufacturing-aware reasoning, we introduce CoT-M: a data organization strategy that infuses Chain-of-Thought (CoT) reasoning with manufacturing processes. CoT-M is built upon InstructIAD and aims to enhance anomaly detection by joint reasoning between product attributes and manufacturing-related information.
CoT-M extends InstructIAD along two axes: (i) product and defect diversity and (ii) manufacturing awareness. The data-generation pipeline (Fig. 2) adapts to the information available for each sample through three complementary modes: (a) Images with captions (Fig. 2, middle): The original attribute-level caption and its MFG are fed into GPT, which interleaves visual details with manufacturing steps to yield a coherent chain-of-thought (CoT). The outcome is an annotated triplet (image, manufacturing process, reasoning trajectory); (b) Images without captions (Fig. 2, bottom): For categories such as “juice bottle” in MVTec-LOCO AD, caption information is absent. We therefore invoke a checklist-style template that emulates a human inspector. Each coarse defect label is cross-referenced with an LLM-generated MFG checklist to pinpoint the faulty visual attribute; the completed checklist itself serves as the CoT explanation. (c) Text-only exemplars (Fig. 2, top): Starting from a normal product caption in InstructIAD, we synthetically augment the product by altering attributes (e.g., colour, component type, quantity) and stochastically injecting defect descriptions. GPT then generates the reasoning trajectory conditioned on the augmented caption and its MFG, producing a purely textual CoT pair. After generating the accompanying reasoning trajectories with the augmented descriptions and manufacturing processes, we manually filter out any hallucinated or erroneous samples. The final output, including an augmented caption, corresponding manufacturing processes, and generated reasoning process forms a new textual CoT answer pair.
@@ 3.5 Expert-guided RoI module
In LMMs, the image is encoded into tokens by a visual encoder such as CLIP [27] and SigLIP [37], thus the encoder resizes the resolution of the input images to a fixed size. To offset the effect of limited resolution, LLaVA-Next [22] has proposed a technique called AnyRes by entering extra large patches of the original image. This technique improves performance and has been further developed in LLaVA-OneVision [16]. However, dividing the whole image into patches as extra patches is memory-consuming and redundant for IAD, as most patches of a defective image do not include defects. Accordingly, we propose the EG‑RoI module (see Fig. 3), which improves the recognition of product attributes and defects. During training, ground‑truth defect regions together with randomly sampled normal regions are supplied to the module. During the evaluation, we can leverage established zero‑shot IAD vision experts—MuSc [18], AnomalyCLIP [40], and April‑GAN [7]—to supply informative suspicious regions to EG‑RoI. Anomaly maps from the vision experts are normalized and binarized; pixels scoring >0.9>0.9 define suspicious regions. Each region is enclosed in a fixed‑size box cropped from the image to preserve resolution. To comply with the language model’s context‑length limit, heavily overlapping boxes are merged, and the total number of regions is capped at four. The resized original image forms the base view, while cropped regions serve as auxiliary patches. Both the base image and the patches are processed by the image encoder; patch tokens are average‑pooled to save memory, flattened, and concatenated with the base‑image tokens. The combined visual tokens are then projected into the textual embedding space and passed as a single sequence to the language model. By supplying these high‑resolution local views, EG‑RoI allows the LMM to conduct a more detailed MFG-driven inspection.
@@ 4 Experiments
[…]
Our experiments mainly focus on how manufacturing processes boost LMMs in IAD tasks through our method, including common results from general LMMs and LMM-based IAD methods and ablation experiments of different industrial contexts.
@@
To preserve the generalization of the base models, we supplement our IAD-related instruction data by sampling 12K pairs from the original fine-tuning datasets of both LLaVA-1.6 and LLaVA-OneVision. Moreover, since LLaVA-1.6 lacks inherent multi-image processing capabilities, we constructed a simple dual-image caption dataset using the COCO subset [21] from the ShareGPT4V dataset [5]. This dataset is exclusively used in the 1-shot setting to provide basic multi-image support. Given the distinct objectives of 0-/1-shot anomaly detection, we offer separate versions of Triad for each setting, with data organized by CoT-M with specific instructions. For the 1-shot version of Triad-llava-1.6, we integrate the zero-shot model with the one-shot model using a Confidence Voting Mechanism (see Sec. E.1) to mitigate its multi-image inability.
@@
For zero-shot evaluation, the instruction follows the same format as the anomaly detection task shown in Fig. 2. For one-shot evaluation, general LMMs that support multi-image directly input the query and a normal reference image with a similar prompt with zero-shot, only the question is replaced with:
@@
LMM-based IAD method AnomalyGPT [10] and Myriad [19] are tested according to their own instructions. Because of their low instruction following ability, we retrieve the keyword yes or no for accuracy calculation. Detailed demonstration can be found in Sec. B.
Baseline In this study, we perform extensive evaluations against several robust baselines, including AnomalyGPT [10] and Myriad [19]. General-purpose LMMs exhibit enhanced reasoning capabilities as a result of being fine-tuned on complex reasoning data spanning diverse domains, including mathematics, chart interpretation, and document analysis. For evaluation, we benchmark our approach against state-of-the-art general-purpose LMMs, including Qwen2-VL [31], MiniCPM-V [34], the LLaVA series [22, 16], and the closed-source GPT-4o [11].
@@ 4.1 Main results
Table 2 shows 1-shot results. Providing a single reference image generally degrades the performance of most general-purpose LMMs due to their limited instruction-following capabilities. Two exceptions, LLaVA-OneVision-ov-7B (91.5% vs. 91.0% in 0-shot) and Qwen2-VL-72B (92.0% vs. 87.3% in 0-shot), still cannot fully utilize manufacturing information for anomaly detection (accuracy down by 5.7% and 1.6%, respectively). Meanwhile, Triad-llava-1.6-7B achieves comparable performance to both AnomalyGPT and Myriad under 1-shot conditions, and significantly outperforms them with manufacturing processes (88.4% vs. 85.4% for Myriad and 80.9% for AnomalyGPT). Leveraging LLaVA-OneVision-ov-8B—a carefully tuned multi-image baseline—Triad-ov-8B achieves 92.9% on MVTec-AD in the 1-shot setting, further improving to 94.1% when incorporating manufacturing processes. These results suggest that even a strong base model can benefit from the Manufacturing processes. Overall, our findings demonstrate that Triad offers robust and flexible anomaly detection across both zero-shot and few-shot scenarios, while the manufacturing-driven IAD paradigm consistently delivers strong performance gains.
@@
[…]
@@
Figure 5: Qualitative evaluation of Triad-llava-ov-7B on MVTec-AD [2]. MuSc [18] is used as vision expert. The related manufacturing process is shown in (a). Three representative examples from the “Pill” class are shown: (a) a normal pill exhibits red speckles on its surface which are a natural outcome of the manufacturing process and could be easily misjudged as defects. (b) Pills with unintended additive variations indicate a defect. (c) a case where Triad correctly predicts a defect despite errors in the expert outputs.
@@
Figure 5 presents three representative cases. In the first example (Fig. 5(a)), a surface discoloration might typically be viewed as a defect (as GPT-4o incorrectly predicts); however, Triad recognizes that the discoloration corresponds to the cranberry additive described in the manufacturing process, leading to a correct classification. Thanks to the EG-RoI module’s ability to capture fine-grained visual cues and the data organization via CoT-M, Triad confidently aligns the observed appearance with the relevant manufacturing steps. In the second example (Fig. 5(b)), a pill containing “yellow speckles” is mistakenly introduced into a production pipeline that should only produce pills with “red speckles.” Triad correctly identifies the inconsistency between the visual observation and the manufacturing procedures. GPT-4o, though aware that red speckles are the standard, fails to accurately match the color of the speckles in question. This underscores the necessity of precise attribute recognition for reasoning. Finally, the third example (Fig. 5(c)) highlights Triad’s resilience to vision expert errors, where it fails to detect a scratch on the surface. Triad still identifies the scratch independently and also notes a yellow discoloration, demonstrating its capacity to rectify external region proposal mistakes by integrating visual analysis with contextual defect reasoning.